# How robust are the parameter constraints extending the $\Lambda$CDM model?

### v1

@@ III.1 Base $\Lambda$CDM parameters
The following parameter to be discussed is the dark matter mass energy density. Figure 2 depicts the marginalized 1D posteriors for $\Omega_{\mathrm{c}}h^{2}$, considering different datasets. As in the case of $\Omega_{\rm b}h^{2}$, the largest departure from the canonical $\Lambda$CDM theory is also reached when either $N_{\rm eff}$ or $A_{\rm lens}$ are free parameters in the fiducial cosmology. Indeed, the degeneracy with $N_{\rm eff}$ can be easily understood in terms of the early Integrated Sachs Wolfe (ISW) effect. This effect is originated from the interaction between CMB photons and the time-dependent gravitational potentials along the line of sight between us and the last scattering surface. Notice that the early ISW effect adds in phase with the primary anisotropy, increasing the height of the first acoustic peaks, with an emphasis on the first one, where the main contribution of the ISW effect lies. In addition, the early ISW effect will be suppressed by the square of the radiation-to-matter ratio $\propto[(1+z_{\rm r})/(1+z_{\rm eq})]^{2}$, i.e. a larger (smaller) matter component will result into a smaller (larger) ISW amplitude due to the larger (smaller) value of $z_{\rm r}$. Conversely, a larger (smaller) amount of radiation – i.e. larger (smaller) $N_{\rm eff}$ – will result into a larger (smaller) ISW effect. This is also the reason for the lower value of $\Omega_{\mathrm{c}}h^{2}$ preferred by ACT observations, due to their slightly different amplitude of the first CMB acoustic peak when compared to Planck data. Concerning $A_{\rm lens}$, notice that there is always a preference for $A_{\rm lens}>1$ when considering Planck or SPT data, see Tab. 10. This implies more CMB lensing and to compensate for that a lower matter density would be required. In the case of SPT, however, we find a value of $A_{\rm lens}<1$, and therefore the value of $\Omega_{\mathrm{c}}h^{2}$ will be higher than in the standard $\Lambda$CDM picture. Finally, the marginalized constraints on the dark matter energy density are very close to those obtained within the minimal scenario, reassuring the stability of the main cosmological parameters, despite the fact that also a time-dependent dark energy equation of state, parameterized via the CPL prescription, or a non-zero curvature, will also have a non-negligible impact on $\Omega_{\rm c}h^{2}$, see Tabs. 12 and 9.
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
The next parameter we shall discuss is $\theta_{\rm MC}$, an approximation to the angular scale of the sound horizon at decoupling, based on some model-dependent analytical fits. In Figs. 3 we depict the one-dimensional probability distributions for this parameter. Table 4 shows the marginalized constraints on $100\theta_{\rm MC}$. The very first thing to notice is that this parameter is very well constrained by Planck CMB measurements, while higher multipole damping-tail CMB probes, such as SPT or ACT, lead to much larger errors. This is due to the fact that Planck observations have an extremely good accuracy around the first acoustic peaks while ACT and SPT are focused on the high multipole region, where it is more difficult to determine the angular position of the peaks because of their smaller amplitudes. Notice as well that low redshift observations from BAO data do not improve much the results. When varying the fiducial cosmology, i.e. allowing for extensions of the minimal $\Lambda$CDM scheme, does not have a very large impact on the mean value and errors of $\theta_{\rm MC}$ except from $N_{\rm eff}$, see Tab. 7. Reference Kosowsky et al. (2002) provides analytical expressions for the sound horizon at decoupling and makes straightforward to understand the large impact of $N_{\rm eff}$ in the mean value and errors of $100\theta_{\rm MC}$. The value of $r_{s}$, the sound horizon at decoupling, is directly linked to the amount of radiation in the universe, governed by $N_{\rm eff}$. Therefore, varying this parameter will have an obvious impact on the value of $100\theta_{\rm MC}$.
We shall comment next on the marginalized limits of the parameter $\tau_{\rm{reio}}$, the reionization optical depth. Table 4 shows two very important points concerning the extraction of this parameter: firstly, it is Planck that obtains the best accuracy on $\tau_{\rm{reio}}$ by means of polarization measurements. Indeed, the errors from either the SPT or ACT datasets are a consequence of the Gaussian prior applied while studying high multipole CMB data from these two experiments, which have no sensitivity on $\tau_{\rm{reio}}$ on their own. Such Gaussian prior is a conservative estimate obtained partially from WMAP and Planck results, see Aiola et al. (2020), which brings twice as large errors as those obtained from Planck CMB measurements. Secondly, notice that low redshift measurements do not basically change the errors on $\tau_{\rm{reio}}$ for Planck, due to their lack of sensitivity to the physical effects induced by this parameter. From what regards the different fiducial cosmologies considered here, the largest impact, as expected, is when the phenomenological parameter $A_{\rm lens}$ is allowed to freely vary, see Tab. 10. Indeed, the mean value of $\tau_{\rm{reio}}$ is smaller than its $\Lambda$CDM counterpart when the lensing amplitude is a free parameter in the analysis. Even if the shift in such a mean value is not significant, it is straightforward to understand. If $A_{\rm lens}>1$, CMB lensing will be stronger. Since $\tau_{\rm{reio}}$ suppresses all acoustic peaks by a constant factor $\exp(-\tau_{\rm{reio}})$, while leaving the power on the largest scales unaffected, a smaller value of $\tau_{\rm reio}$, leading to a smaller smearing of the small-scale features in the power spectrum, will be required to compensate for such an effect. Nevertheless, the marginalized constraints on the reionization optical depth are extremely stable and pretty robust.
The next parameter we discuss is the primordial power spectrum scalar spectral index, $n_{\rm s}$. Table 4 illustrates the fact that low redshift measurements do not imply a major improvement when combined with Planck CMB measurements. Instead, they improve the accuracy in the extraction of $n_{\rm s}$ when added to SPT or ACT observations. Figure 5 illustrate the one-dimensional probability distribution for the scalar spectral index parameter for different data sets and fiducial cosmologies. The largest departure takes place when the number of relativistic degrees of freedom, $N_{\rm{eff}}$, is also a free parameter in the fiducial cosmology, see Tab. 7. The effect is straightforward to understand: values $N_{\rm{eff}}<3$ would imply less damping in the CMB high multipole region. This effect can be compensated by a lower value of $n_{\rm s}$, that will decrease the slope of the angular power spectrum, lowering the right side with respect to the left side. Conversely values $N_{\rm{eff}}>3$ requires the mean value of $n_{\rm s}$ to increase in order to compensate for the larger damping induced by a larger dark radiation component. Nevertheless, the marginalized errors are pretty robust and close to their values within the minimal $\Lambda$CDM scenario.
Refer to captionRefer to caption
@@
Refer to captionRefer to caption

### v2

@@ III.2 Cold Dark Matter Energy Density $\Omega_{\mathrm{c}}h^{2}$
The next parameter to be discussed is the dark matter mass-energy density. As usual, we start by scrutinizing the main correlations with non-standard cosmological parameters beyond the $\Lambda$CDM cosmology. We then derive marginalized constraints on this quantity, accounting for a wide range of minimal extensions across all six combinations of data.
III.2.1 Main correlations with beyond $\Lambda$CDM parameters
When it comes to correlations among parameters, the situation for the cold dark matter energy density turns out to be very similar to that of $\Omega_{\rm b}h^{2}$. In this case, the largest departure from the canonical $\Lambda$CDM theory is also observed when either $N_{\rm eff}$ or $A_{\rm lens}$ are free parameters in the fiducial cosmology. For all the combinations of data explored in this study, the one-dimensional posterior distribution functions and the two-dimensional correlations between $\Omega_{\rm c}h^{2}$, $N_{\rm eff}$, and $A_{\rm lens}$ are shown in the second row from the top in Fig. 1. The left panel displays the correlation between $\Omega_{\rm c}h^{2}$ and $A_{\rm lens}$, while the right panel shows the correlation between $\Omega_{\rm c}h^{2}$ and $N_{\rm eff}$.
@@ III.2.1 Main correlations with beyond $\Lambda$CDM parameters
The degeneracy between $\Omega_{\rm c}h^{2}$ and $N_{\rm eff}$ can be easily understood in terms of the early Integrated Sachs Wolfe (ISW) effect. This effect is originated from the interaction between CMB photons and the time-dependent gravitational potentials along the line of sight between us and the last scattering surface. Notice that the early ISW effect adds in phase with the primary anisotropy, increasing the height of the first acoustic peaks, with an emphasis on the first one, where the main contribution of the ISW effect lies. In addition, the early ISW effect will be suppressed by the square of the radiation-to-matter ratio $\propto[(1+z_{\rm r})/(1+z_{\rm eq})]^{2}$, i.e. a larger (smaller) matter component will result into a smaller (larger) ISW amplitude due to the larger (smaller) value of $z_{\rm r}$. Conversely, a larger (smaller) amount of radiation – i.e. larger (smaller) $N_{\rm eff}$ – will result into a larger (smaller) ISW effect. This is also the reason for the lower value of $\Omega_{\mathrm{c}}h^{2}$ preferred by ACT observations, due to their slightly different amplitude of the first CMB acoustic peak when compared to Planck data.
Concerning $A_{\rm lens}$, notice that there is always a preference for $A_{\rm lens}>1$ when considering Planck or SPT data, see Tab. 9. This implies more CMB lensing and to compensate for that a lower matter density would be required. In the case of SPT, however, we find a value of $A_{\rm lens}<1$ and therefore the value of $\Omega_{\mathrm{c}}h^{2}$ will be higher than in the standard $\Lambda$CDM picture (see also Fig. 1).
III.2.2 Marginalized Constraints
Figure 3 depicts the displacement in the one-dimensional posterior distribution functions for $\Omega_{\mathrm{c}}h^{2}$ resulting from assuming different cosmologies, along with the marginalized 1D posteriors highlighted in bold, considering various datasets. Overall, the marginalized constraints on the dark matter energy density are very close to those obtained within the minimal $\Lambda$CDM scenario, reassuring the stability of the main cosmological parameters. The constraints obtained from the different CMB datasets agree quite well with each other, both when including and excluding low-redshift observations. However, we note that including low-redshift data is critical for achieving strong constraining power in the marginalized constraints. In contrast to the constraints on the baryon density discussed in the previous section (where the constraining power was relatively insensitive to the inclusion of low-redshift data), here incorporating local universe information can lead to an improvement in the resulting constraining power of more than a factor of two. This is due to the fact that a time-dependent dark energy equation of state, parameterized via the CPL prescription, or a non-zero curvature will also have a significant impact on $\Omega_{\rm c}h^{2}$ (see Tabs. 11 and 8). These parameters, which affect the late-time evolution of the universe, are almost unconstrained without distance observations but are well constrained when we add BAO and SN data.
@@ III.2.2 Marginalized Constraints
III.3 The Angular scale of the Sound Horizon, $\theta_{\rm MC}$
The next parameter we shall discuss is $\theta_{\rm MC}$, an approximation to the angular scale of the sound horizon at decoupling, based on some model-dependent analytical fits.
III.3.1 Main correlations with beyond $\Lambda$CDM parameters
Looking at the constraints for the $\theta_{\rm MC}$ parameter in Tabs. 411 for the different models, the first thing to notice is that $\theta_{\rm MC}$ is very well constrained by Planck CMB measurements. In contrast, higher multipole damping-tail CMB probes like SPT and ACT result in much larger errors. This is because Planck data has extremely high accuracy around the first acoustic peaks, while ACT and SPT focus on the high multipole region, where determining the angular position of the peaks is more challenging due to their smaller amplitudes. Additionally, low-redshift observations from BAO data do not significantly improve the constraints on $\theta_{\rm MC}$.
When varying the fiducial cosmology by allowing extensions beyond the minimal $\Lambda$CDM model, the impact on the mean value and errors of $\theta_{\rm MC}$ is generally small. The only notable exception is when $N_{\rm eff}$ is varied, as shown in Tab. 6. In Fig. 1 (third row, right panel), we plot the correlations between $N_{\rm eff}$ and $\theta_{\rm MC}$ for the different CMB experiments, both individually and in combination with low-redshift data. As seen in the figure, there is a quite strong negative correlation, especially in the ACT and SPT-based constraints. Increasing (or decreasing) the effective number of relativistic particles decreases (or increases) the angular scale of the sound horizon.
This correlation can be understood by referring to the analytical expressions for the sound horizon at decoupling, as discussed in Ref. Kosowsky et al. (2002). Specifically, $r_{s}$, the sound horizon at decoupling, is directly linked to the amount of radiation in the universe, which is governed by $N_{\rm eff}$: increasing the radiation energy density decreases the size of the sound horizon. In turn, $100\theta_{\rm MC}$ is essentially the ratio of the sound horizon to the angular diameter distance to the CMB, which contains information about the universe’s expansion history after recombination. If we don’t change the late-time cosmology but only vary $N_{\rm eff}$ (or use low-redshift observations that constrain the post-recombination expansion to be very close to the Planck $\Lambda$CDM cosmology), this affects the numerator of this ratio without significantly changing the denominator. This has a clear impact on the inferred value of $100\theta_{\rm MC}$. This explains the relatively large effect that $N_{\rm eff}$ has on both the mean value and errors of $100\theta_{\rm MC}$, as shown in Fig. 1.
III.3.2 Marginalized Constraints
As usual, in Fig. 4, we depict the one-dimensional probability distributions for this parameter obtained from different combinations of datasets within the various fiducial cosmologies analyzed in this work. The bold lines represent the marginalized one-dimensional posterior probability distribution functions. The first noticeable point is that the displacement between the marginalized posterior and those obtained from different fiducial cosmologies is smaller than for other parameters, with the only exception being the results from varying $N_{\rm eff}$, as mentioned in the previous subsection.
The numerical results for $100\theta_{\rm MC}$ are summarized in Table 3, both at the 68% and 95% CL. These results confirm that Planck-based constraints determine this parameter with a precision more than twice as accurate as those based on ACT and SPT. This is due to the lack of data from ACT and SPT around the first acoustic peak, which is crucial for precisely measuring the angular scale of the sound horizon.
Additionally, from Figs. 4, we note that ACT-based constraints are shifted towards larger values of $\theta_{\rm MC}$ compared to Planck, while SPT-based constraints are shifted towards smaller values. This results in a mild ( $2\sigma$) difference in the constraints on this parameter from different CMB probes. The situation remains similar when considering constraints from the combination of CMB experiments and low-redshift data. However, in these cases, we observe a reduction in the displacement of the one-dimensional probability distributions (see the right panel of Fig. 4) and a decrease in uncertainties for ACT and SPT-based marginalized constraints. On the other hand, the improvements for Planck-based constraints are minimal. This is because Planck data typically constrain this parameter well, even in extended cosmologies, due to the availability of data around the first peak. In contrast, ACT and SPT constraints allow more flexibility in shifting $\theta_{\rm MC}$ when studying fiducial cosmologies that modify the late-time expansion history. In these latter cases, the angular diameter distance to the CMB is well constrained only when low-redshift data are included, resulting in better-determined values for $\theta_{\rm MC}$.
We conclude this subsection with an important remark on the utility of our marginalized constraints. Specifically, the value of the angular scale of the sound horizon measured by Planck under the assumption of a fiducial $\Lambda$CDM cosmology is often used as a distance prior to compress CMB likelihoods and constrain extensions to the standard cosmological model. Essentially, the angular scale of the sound horizon can be viewed as an additional BAO data point at recombination (i.e., at $z\sim 1090$) to be considered when testing new models. This parameter is commonly used because, as our analysis partially confirmed, $100\theta_{\rm MC}$ is considered a relatively model-independent geometric measurement characterizing the CMB angular power spectra (being closely tied to the position of the first acoustic peak). However, the marginalized result we provide in Table 3 already accounts for potential effects arising from varying a wide range of fiducial cosmologies considered here, thereby reducing the model dependence of our estimate. Therefore, this result can and should be used in such analyses instead of the most economical $\Lambda$CDM one, replacing the value obtained from assuming a single baseline cosmology only.
@@ III.3.2 Marginalized Constraints
III.4 Reionization optical depth, $\tau_{\rm{reio}}$
We shall comment next on the correlations and marginalized limits of the parameter $\tau_{\rm{reio}}$, the reionization optical depth.
III.4.1 Main correlations with beyond $\Lambda$CDM parameters
Constraints on $\tau_{\rm{reio}}$ based on Planck (and eventually ACT) CMB measurements, as well as its correlation with various beyond-$\Lambda$CDM parameters, have recently undergone extensive testing. Specifically, in Ref. Giarè et al. (2023a), some of us analyzed the reionization optical depth in relation to large-scale temperature and polarization measurements, investigating the possibility that anomalies in these parts of the datasets could lead to artificially low values of $\tau_{\rm{reio}}$, thereby driving several minor or mild anomalies in the Planck constraints. For more details in this direction, we refer to the aforementioned work. Here, we confirm that Planck’s accuracy on $\tau_{\rm{reio}}$ is largely driven by large-scale temperature and polarization measurements. In contrast, the errors from either the SPT or ACT datasets arise from the Gaussian prior applied when studying high-multipole CMB data from these two experiments, which, on their own, have no sensitivity to $\tau_{\rm{reio}}$. It’s worth noting that this Gaussian prior is a conservative estimate, partially derived from WMAP and Planck results (see Ref. Aiola et al. (2020)), and it results in errors twice as large as those obtained from Planck CMB measurements. [^4] In Ref. Giarè et al. (2023a), some of us argued that, in principle, by combining low-redshift data with small-scale ACT CMB measurements, it is possible to constrain $\tau_{\rm{reio}}$ without adopting such a prior, thereby obtaining a Planck-independent estimate of this parameter. The resulting constraints tend to favor slightly larger values of the optical reionization depth, showing a $\sim 1.8\sigma$ tension with the Planck result. That being said, these constraints are significantly weaker in extended cosmological models, as demonstrated in the same work. This is the primary reason we chose to adopt the prior in our study. Additionally, it should be noted that low-redshift measurements do not significantly alter the errors on $\tau_{\rm{reio}}$ obtained by Planck, due to their lack of sensitivity to the physical effects induced by this parameter.
Regarding the different fiducial cosmologies considered in this work, the most significant impact occurs, as expected, when the phenomenological parameter $A_{\rm lens}$ is allowed to vary freely (see Tab. 9 for numerical results and the third row, left panel of Fig. 1 for the one-dimensional posterior distributions and two-dimensional correlations between $\tau_{\rm{reio}}$ and $A_{\rm lens}$). As seen in the figure, due to this correlation, the mean value of $\tau_{\rm{reio}}$ is lower than its $\Lambda$CDM counterpart when the lensing amplitude is a free parameter in the analysis. Although the shift in the mean value is not significant, it is straightforward to understand: if $A_{\rm lens}>1$, CMB lensing will be stronger. Since $\tau_{\rm{reio}}$ suppresses all acoustic peaks by a constant factor $\exp(-\tau_{\rm{reio}})$ while leaving the power on the largest scales unaffected, a smaller value of $\tau_{\rm{reio}}$ – leading to reduced smearing of the small-scale features in the power spectrum – will be required to compensate for this effect.
III.4.2 Marginalized Constraints
The marginalized constraints on $\tau_{\rm{reio}}$ are shown in Table 3 and Fig. 5. In this case, we note that the constraints on the reionization optical depth are highly stable and robust. For the ACT and SPT-based results, this stability is primarily due to the inclusion of the Gaussian prior. Minor shifts in the one-dimensional posterior distribution functions obtained within different extended cosmological models are observed, indicating that even with a prior on this parameter, slight displacements can arise when varying other quantities simultaneously. Our marginalized distributions (highlighted in bold in the figure) account for these minor shifts, weighting them by the Bayes factor of the different models.
@@ III.4.2 Marginalized Constraints
III.5 Amplitude of the primordial power spectrum $A_{\rm s}$
Next, we shall focus on the amplitude of the primordial power spectrum $A_{\rm s}$.
III.5.1 Main correlations with beyond $\Lambda$CDM parameters
When it comes to correlations between $\tau_{\rm{reio}}$ and beyond-$\Lambda$CDM parameters, we find very similar results to those discussed earlier for $\tau_{\rm{reio}}$. This is primarily due to the strong degeneracy between the amplitude of the primordial power spectrum, $A_{\rm s}$, and the optical reionization depth, $\tau_{\rm{reio}}$. A well-known effect of reionization, aside from enhancing the CMB polarization spectrum at large angular scales, is the suppression of temperature anisotropies at smaller scales. This suppression can be offset by increasing $A_{\rm s}$, which controls the overall normalization of the CMB power spectrum. Therefore, it is not surprising that the same parameters that correlate with $\tau_{\rm{reio}}$ also affect the amplitude of the primordial spectrum.
As shown in Tab. 9, the largest impact on the determination of $A_{\rm s}$, particularly on its error, occurs when the lensing amplitude, $A_{\rm lens}$, is a free parameter. The one-dimensional posterior distribution functions and two-dimensional correlations for these parameters are presented in the bottom-left panel of Fig. 1, and they exhibit a behavior similar to that described for the optical depth. Specifically, to counterbalance the effects of considering more (or less) lensing than expected in the standard $\Lambda$CDM model, one would need to decrease (or increase) the primordial amplitude. As seen in the figure and in Tab. 9, low-redshift data do not significantly improve the constraining power in this case.
III.5.2 Marginalized Constraints
The marginalized constraints for $A_{\rm s}$ are summarized in Tab. 3 and visualized in the two panels of Fig. 6, where we also show the displacements in the one-dimensional probability distribution for this parameter across different datasets and fiducial cosmologies, together with the resulting marginalized one-dimensional posteriors in bold. Overall, the central values agree quite well across the various combinations of data, typically within one standard deviation. However, for CMB experiments other than Planck, the error budget is nearly doubled. This discrepancy in precision between Planck and Planck-independent experiments is somewhat mitigated by the inclusion of low-redshift data, although the uncertainty remains significantly larger than that for Planck combined with low-z data. In conclusion, low-redshift observations provide some improvement in reducing uncertainties in Planck-independent constraints, but Planck continues to offer superior precision in determining $A_{\rm s}$ across all cosmological models and datasets.
Figure 8: Marginalized 1D and 2D posteriors involving the most relevant derived cosmological parameters ($H_{0}$, $\Omega_{m}$, $\sigma_{8}$) and other beyond-$\Lambda$CDM parameters that exhibit the most significant correlations and degeneracy lines with the former, across all the CMB and CMB+low-$z$ datasets analyzed in this work.
@@ III.5.2 Marginalized Constraints
III.6 Primordial scalar spectral index $n_{\rm s}$
The next parameter we discuss is the primordial power spectrum scalar spectral index, $n_{\rm s}$.
III.6.1 Main correlations with beyond $\Lambda$CDM parameters
Constraints on the value of the spectral index of primordial fluctuations have recently been a matter of intense debate in the literature for several reasons. First and foremost, as highlighted by our extensive reanalysis, mild to moderate differences in the cosmological parameters inferred by ACT and Planck have emerged, as reported by several independent groups, including some of us, in Refs. Lin and Ishak (2021); Forconi et al. (2021); Handley and Lemos (2021); La Posta et al. (2023); Di Valentino et al. (2022a, b); Giarè et al. (2022); Calderón et al. (2023); Giarè (2023). These discrepancies primarily involve the spectral index of primordial fluctuations, $n_{s}$, or inflationary parameters closely related to $n_{s}$ (see e.g., Refs. Forconi et al. (2021); Giarè et al. (2022); Di Valentino et al. (2022b) for further discussion). Unexpectedly, small-scale CMB measurements from the Data Release 4 of the Atacama Cosmology Telescope (ACT) show agreement with a scale-invariant Harrison-Zel’dovich spectrum ($n_{s}\sim 1$), introducing a tension with Planck results at 99.3% CL  Giarè et al. (2022). If observational systematics are set aside, and the differences between ACT and Planck are considered as genuine, the analysis of small-scale CMB data leads to distinct predictions for inflationary models Giarè et al. (2023b). Secondly, in Ref. Giarè (2024), one of us extensively investigated how Planck’s constraints on $n_{s}$ evolve in beyond-$\Lambda$CDM theoretical models aimed at resolving recent cosmological tensions (which we will discuss later in more detail). This analysis pointed out that deviations from a baseline $\Lambda$CDM cosmology at early times (i.e., before recombination) can significantly affect the results on the spectral index.
Unsurprisingly, our analysis confirms that the largest impact on the constraints on $n_{s}$ occurs when the number of relativistic degrees of freedom, $N_{\rm{eff}}$, is a free parameter in the fiducial cosmology (see Tab. 6). The correlations between $n_{s}$ and $N_{\rm{eff}}$ are shown in the bottom-right panel of Fig. 1, where all datasets reveal a strong positive correlation between these two parameters. This effect is straightforward to understand: values of $N_{\rm{eff}}<3$ imply less damping in the high-multipole region of the CMB. A lower $N_{\rm{eff}}$ can be compensated by a lower $n_{s}$, which decreases the slope of the angular power spectrum, reducing the right side relative to the left one. Conversely, values of $N_{\rm{eff}}>3$ require the mean value of $n_{s}$ to increase to compensate for the larger damping induced by a larger dark radiation component. The figure also reveals that varying these two parameters simultaneously in the cosmological model tends to reduce the preference for $n_{s}\sim 1$ observed in ACT-based results. However, this comes at the cost of shifting $N_{\rm{eff}}$ towards slightly lower values than those expected within the standard model of cosmology and particle physics (see Ref. Di Valentino et al. (2022a) for further discussion).
III.6.2 Marginalized Constraints
The marginalized constraints on the spectral index are summarized in Tab. 3 at 68% and 95% CL. As shown in Fig. 7, the scatter in the one-dimensional marginalized posteriors obtained from different fiducial cosmologies is quite significant. This underscores that the constraints on this parameter depend non-trivially on the overall cosmology, as previously discussed. Notably, the figure confirms an overall tendency for the posterior distributions based on ACT and SPT CMB measurements to prefer slightly larger values than those inferred from Planck CMB data, consistent with the references outlined earlier. In fact, the marginalized limit on $n_{s}$ from ACT data reported in Table 3 aligns with $n_{s}\sim 1$ and is significantly shifted compared to Planck results. SPT constraints are generally in line with Planck but have such large uncertainties that they remain consistent with ACT results as well. Among the three CMB experiments, the marginalized posterior distribution functions that most strongly rule out $n_{s}\sim 1$ are those based on Planck data. Even in this case, the error bars on the spectral index are significantly larger than those obtained within a baseline $\Lambda$CDM cosmology. Last but not least, the same figure and table illustrate that low-redshift measurements do not lead to substantial improvements when combined with Planck CMB data. Instead, they slightly enhance the accuracy of $n_{s}$ extraction when added to SPT or ACT observations.
III.7 Hubble Constant, $H_{0}$
We have discussed so far the main six $\Lambda$CDM parameters. Now we will focus on the derived ones, starting with the Hubble constant, $H_{0}$.